### v3

@@ Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (Bright) to support AI-based all-weather disaster response. To the best of our knowledge, Bright is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of human-made disasters across 14 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in this dataset, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In the experiments, we have tested seven advanced AI models trained with Bright to validate the transferability and robustness. Beyond that, it also serves as a challenging benchmark for a variety of tasks under disaster scenarios, including unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM). The experimental results serve as baselines to inspire future research and model development. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as new disaster data becomes available. Bright also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven methods in support of people in disaster-affected areas.
@@
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods have affected 4.4 billion people and caused approximately 1.3 million deaths. These disasters have also resulted in economic losses of US\$2,647 billion in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions to minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, preventing efficient on-site assessments. Earth observation (EO), also known as remote sensing, provides a safe and efficient way to obtain information on building damage in disaster areas because of its wide field of view and contactless operation.
@@
The EO technologies commonly used for assessing building damage after disasters are optical and synthetic aperture radar (SAR). Optical imagery is a primary source for building damage assessment because of its intuitive and easy-to-interpret nature. For example, moderate-resolution optical EO data from the Landsat series and Sentinel-2 have been used to assess building damage (Yusuf et al., 2001; Fan et al., 2019; Sandhini Putri et al., 2022). Landsat and Sentinel-2 data are limited in spatial resolution and only provide broad approximations of affected areas, which lack accuracy for specific buildings crucial for timely rescue. The new generation of very-high-resolution (VHR) optical sensors, such as IKONOS and WorldView, provide EO data with spatial resolutions at a meter and a sub-meter level, enabling finer assessments at the level of individual buildings (Freire et al., 2014). These data have been used successfully for the assessment of building damage (Yamazaki and Matsuoka, 2007; Tong et al., 2012; Freire et al., 2014).
@@
While accurate building damage maps can be obtained by human-expert visual interpretation of optical images, this process is time-consuming and labor-intensive for rapid large-scale assessments. In addition, it requires trained professionals. Therefore, recent literature has focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, machine learning (ML) and deep learning (DL) techniques have significantly improved the efficiency and accuracy of building damage assessment. Earlier work was toward a single disaster event with labels annotated for a particular disaster area to train a model, which is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) of different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Chen et al., 2022a; Shen et al., 2022; Kaur et al., 2023; Guo et al., 2024; Wang et al., 2024; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damages of buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
@@
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1c is not affected by smoke and clearly shows the buildings damaged by the wildfire. Due to the advantages of SAR imagery, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate accurate building damage mapping compared to optical imagery (Adriano et al., 2021; Xia et al., 2025). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010; Adriano et al., 2021). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This is a significant challenge that remains and needs to be addressed.
@@
ABCD (Fujita et al., 2017) Yes Optical EO 0.4 1 Tsunami N/A Image-level
(Nguyen et al., 2017) Yes Images on social media N/A 4 3 natural disasters N/A Image-level
(Cheng et al., 2021) Yes UAV image N/A 1 Hurricane 1,802 Image-level
(Xue et al., 2024) Yes Street-view image N/A 1 Hurricane 2,468 Image-level
FloodNet (Rahnemoonfar et al., 2021) Yes UAV image N/A 1 Flood 6,675 Pixel-level
RescueNet (Rahnemoonfar et al., 2023) Yes UAV image N/A 1 Hurricane 10,903 Pixel-level
Ida-BD (Kaur et al., 2023) No Optical EO 0.5 1 Hurricane 18,083 Pixel-level
xBD (Gupta et al., 2019) Yes Optical EO <0.8 15 6 natural disasters >700,000 Pixel-level
QQB (Sun et al., 2024b) Yes Optical and SAR EO <1 1 Earthquake 4,029 Pixel-level
BDD (Adriano et al., 2021) No Optical and SAR EO 1.2-3.3 9 3 natural disasters 123,453 Pixel-level
Bright Yes Optical and SAR EO 0.3-1 14 5 natural disasters 2 human-made disasters 384,596 Pixel-level
@@ 0.1 Comparison with existing datasets
The comparison between Bright and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to limited available disaster events with corresponding open-source EO data and annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. However, the xBD includes only optical EO data. It does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models. The dataset most similar to Bright is the BDD proposed by Adriano et al. (2021). The main differences between BDD and Bright datasets are: (1) Bright covers more disaster events and building instances, including both natural and human-made disasters. (2) Bright has higher spatial resolution; while the highest resolution of SAR images in BDD is 1.2 meters, Bright provides finer detail with spatial resolution from 0.3 meters to 1 meter to enable the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, Bright is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not concerned with building damage assessment.
@@ 0.2 Main contribution
The contributions of this paper are twofold:
@@
    We present Bright, the first multimodal building damage dataset with sub-meter spatial resolution, which is publicly available to the community. Bright employs a combination of pre-event optical imagery and post-event SAR imagery with various disaster events and rich geographic diversity to support the study of AI-based multimodal building damage mapping, especially in developing countries.
@@
    We evaluate a suite of contemporary models on Bright to establish robust baselines. Beyond supervised deep learning, Bright can support a wide range of AI-based methods. It enables research in UDA, SSL, UMCD, and UMIM, among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with source code and pretrained model weights, are publicly released to provide strong baselines and accelerate future developments for disaster response in the community.
1 Study areas and disaster events
Figure 2: Geographic distribution of disaster events present in the Bright.
We selected 14 disaster events across the globe for Bright, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts.
Table 2: Summary of the basic information of the Bright dataset, with disaster events listed in chronological order.
@@ 1 Study areas and disaster events
Beirut, Lebanon Explosion (EP) 04 Aug. 2020 1 Maxar &\&& Capella 133 25,496
Bata, Equatorial Guinea Explosion (EP) 07 Mar. 2021 0.5 Maxar &\&& Capella 107 8,893
Goma, DR Congo Volcano eruption (VE) 22 May 2021 0.33 Maxar &\&& Capella 123 18,741
Les Cayes, Haiti Earthquake (EQ) 14 Aug. 2021 0.48 Maxar &\&& Capella 73 18,918
La Palma, Spain Volcano Eruption (VE) 19 Sept. 2021 - 13 Dec. 2021 0.3-0.35 IGN (Spain) &\&& Capella 933 30,239
Boulder, USA Wildfire (WF) 30 Dec. 2021 - 01 Jan., 2022 0.6 NAIP &\&& Capella 77 8,365
Ukraine Armed conflict (AC) 22 Mar. 2022 - 21 Sept. 2022 0.6 Google Earth &\&& Capella 513 56,770
Turkey Earthquake (EQ) 06 Feb. 2023 0.30-0.35 Maxar &\&& Capella &\&& Umbra 1,114 135,033
Kyaukpyu, Myanmar Cyclone (CC) 14 May 2023 0.6 Google Earth &\&& Capella 126 8,052
Maui, Hawaii, USA Wildfire (WF) 08 Aug. 2023 - 09 Aug. 2023 0.6 NOAA &\&& Capella 65 3,995
Morocco Earthquake (EQ) 08 Sept. 2023 0.35-0.4 Maxar &\&& Capella 567 6,269
Derna, Libya Flood (FL) 10 Sept. 2023 0.35 Maxar &\&& Capella 124 1,0979
Acapulco, Mexico Hurricane (HC) 25 Oct. 2023 0.35-0.8 Google Earth &\&& Capella 212 18,437
Noto, Japan Earthquake (EQ) 01 Jan. 2024 0.5 GSI &\&& Umbra 79 8,153

### v4

@@ Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability crucial in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on developing artificial intelligence (AI) models to accurately map unseen disaster events, mostly using optical EO data. These solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly combining optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the lack of suitable benchmark datasets has constrained the development of robust multimodal AI models. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (Bright) to support AI-based all-weather disaster response. To the best of our knowledge, Bright is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of human-made disasters across 14 regions worldwide, focusing on developing countries where external assistance is most needed. The dataset’s optical and SAR images with spatial resolutions between 0.3 and 1 meters provide detailed representations of individual buildings, making it ideal for precise damage assessment. We train seven advanced AI models on Bright to validate transferability and robustness. Beyond that, it also serves as a challenging benchmark for a variety of tasks in real-world disaster scenarios, including unsupervised domain adaptation, semi-supervised learning, unsupervised multimodal change detection, and unsupervised multimodal image matching. The experimental results serve as baselines to inspire future research and model development. The dataset (Chen et al., 2025), along with the code and pretrained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as and when a new disaster data is available. Bright also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven methods in support of people in disaster-affected areas.
@@
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods affected approximately 4.4 billion people and caused 1.3 million deaths. These disasters have also resulted in economic losses of 2,647 billion United States dollars (USD) in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions that minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster can be difficult and dangerous, especially when transportation and communication systems are disrupted, making efficient on-site assessments challenging. Earth observation (EO) provides a safe and efficient way to obtain information on building damage in disaster areas due to its wide field of view and contactless operation.
The EO technologies commonly used for assessing building damage after disasters are optical and synthetic aperture radar (SAR). Optical imagery is a primary source for building damage assessment because of its intuitive and easy-to-interpret nature. For example, moderate-resolution optical data from the Landsat series and Sentinel-2 have been used to assess building damage (Yusuf et al., 2001; Fan et al., 2019; Sandhini Putri et al., 2022). Landsat and Sentinel-2 data are limited in spatial resolution and only provide broad approximations of affected areas, which lack precision for specific buildings, crucial for timely rescue. The new generation of very high-resolution (VHR) optical sensors, such as IKONOS and WorldView, provides EO data with spatial resolutions of a meter or less, enabling finer assessments at the level of individual buildings (Freire et al., 2014). These data have been used successfully in building damage assessment (Yamazaki and Matsuoka, 2007; Tong et al., 2012; Freire et al., 2014).
@@
While accurate building damage maps can be obtained by visual interpretation of optical images by human experts, this process is time-consuming and labor-intensive for large-scale rapid assessments. In addition, it requires trained professionals. Therefore, recent studies have focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, machine learning (ML) and deep learning (DL) techniques have significantly improved efficiency and accuracy in building damage assessment. Earlier work focused on a single disaster event with labels annotated for a specific disaster area to train a model. This model is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) containing different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Chen et al., 2022a; Shen et al., 2022; Kaur et al., 2023; Guo et al., 2024; Wang et al., 2024; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damage to buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1-(b) does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1-(c) is not affected by smoke and clearly shows the buildings damaged by the wildfire.
Due to the advantages of SAR imagery, various SAR-based methods have been proposed for building damage assessment. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown.
@@
The inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate the accurate mapping of building damage compared to optical imagery (Adriano et al., 2021; Xia et al., 2025). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010; Adriano et al., 2021). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality, large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This remains a significant challenge that needs to be addressed.
@@
ABCD (Fujita et al., 2017) $\checkmark$ Optical EO 0.4 1 Tsunami N/A Image-level
(Nguyen et al., 2017) $\checkmark$ Images on social media N/A 4 3 natural disasters N/A Image-level
(Cheng et al., 2021) $\checkmark$ Optical EO N/A 1 Hurricane 1,802 Image-level
(Xue et al., 2024) $\checkmark$ Street-view image N/A 1 Hurricane 2,468 Image-level
FloodNet (Rahnemoonfar et al., 2021) $\checkmark$ Optical EO N/A 1 Flood 6,675 Pixel-level
RescueNet (Rahnemoonfar et al., 2023) $\checkmark$ Optical EO N/A 1 Hurricane 10,903 Pixel-level
Ida-BD (Kaur et al., 2023) $\times$ Optical EO 0.5 1 Hurricane 18,083 Pixel-level
CRASAR-U-DROIDs $\checkmark$ Optical EO 0.02-0.12 10 4 natural disasters 21,716 Pixel-level
(Manzini et al., 2024) 1 man-made disaster
Noto-BDA-MV (Vescovo et al., 2025) $\checkmark$ Optical EO N/A 1 Earthquake 140,208 Pixel-level
xBD (Gupta et al., 2019) $\checkmark$ Optical EO <0.8 15 6 natural disasters >700,000 Pixel-level
QQB (Sun et al., 2024b) $\checkmark$ Optical and SAR EO <1 1 Earthquake 4,029 Pixel-level
BDD (Adriano et al., 2021) $\times$ Optical and SAR EO 1.2-3.3 9 3 natural disasters 123,453 Pixel-level
Bright $\checkmark$ Optical and SAR EO 0.3-1 14 5 natural disasters 384,596 Pixel-level
2 man-made disasters
@@ 0.1 Comparison with existing datasets
The comparison between Bright and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to the limited availability of disaster events with corresponding open-source EO data and annotation efforts (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. However, the xBD includes only optical EO data. It does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models.
The dataset most similar to Bright is the BDD proposed by Adriano et al. (2021). The main differences between BDD and Bright datasets are: (1) Bright covers more disaster events and building instances, including both natural and human-made disasters. (2) Bright has higher spatial resolution in both optical and SAR images. Whereas the highest resolution of SAR images in BDD is 1.2 meters, Bright provides finer detail with spatial resolutions ranging from 0.3 meters to 1 meter, enabling the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, Bright is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not related to building damage assessment.
Figure 2: Geographic distribution of disaster events present in Bright.
@@ 0.2 Main contribution
The contributions of this paper are threefold:
@@
    We present Bright, the first multimodal building damage dataset with sub-meter spatial resolution, which is publicly available to the community. Bright employs a combination of pre-event optical imagery and post-event SAR imagery, with various disaster events and rich geographic diversity, to support the study of AI-based multimodal building damage mapping, especially in developing countries.
@@
    We evaluate a suite of contemporary models on Bright to establish robust baselines. Beyond supervised deep learning, Bright can support a wide range of AI-based methods. It enables research in unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM), among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with the source code and pretrained weights, are publicly released to provide strong baselines and accelerate future developments in disaster response in the community.
  3. 3.
    We provide an in-depth analysis that uncovers key challenges and mechanisms of the multimodal building damage assessment task. Through carefully designed experiments, we reveal the difficulties of cross-event generalization, investigate the role of pre-event optical data in aiding damage classification, and quantify the performance gaps between different post-event modalities. These findings offer valuable insights for the development of more robust and practical models for disaster response.
Table 2: Summary of basic information of the Bright dataset with disaster events listed in chronological order. GSI refers to the Geospatial Information Authority of Japan, and IGN refers to the Instituto Geográfico Nacional (National Geographic Institute) of Spain.
@@
Beirut, Lebanon Explosion (EP) 04 Aug. 2020 1 Maxar &\& Capella 133 25,496
Bata, Equatorial Guinea Explosion (EP) 07 Mar. 2021 0.5 Maxar &\& Capella 107 8,893
Goma, DR Congo Volcano eruption (VE) 22 May 2021 0.33 Maxar &\& Capella 123 18,741
Les Cayes, Haiti Earthquake (EQ) 14 Aug. 2021 0.48 Maxar &\& Capella 73 18,918
La Palma, Spain Volcano Eruption (VE) 19 Sept. 2021 - 13 Dec. 2021 0.3-0.35 IGN (Spain) &\& Capella 933 30,239
Boulder, USA Wildfire (WF) 30 Dec. 2021 - 01 Jan., 2022 0.6 NAIP &\& Capella 77 8,365
Ukraine Armed conflict (AC) 22 Mar. 2022 - 21 Sept. 2022 0.6 Google Earth &\& Capella 513 56,770
Turkey Earthquake (EQ) 06 Feb. 2023 0.30-0.35 Maxar &\& Capella &\& Umbra 1,114 135,033
Kyaukpyu, Myanmar Cyclone (CC) 14 May 2023 0.6 Google Earth &\& Capella 126 8,052
Maui, Hawaii, USA Wildfire (WF) 08 Aug. 2023 - 09 Aug. 2023 0.6 NOAA &\& Capella 65 3,995
Morocco Earthquake (EQ) 08 Sept. 2023 0.35-0.4 Maxar &\& Capella 567 6,269
Derna, Libya Flood (FL) 10 Sept. 2023 0.35 Maxar &\& Capella 124 1,0979
Acapulco, Mexico Hurricane (HC) 25 Oct. 2023 0.35-0.8 Google Earth &\& Capella 212 18,437
Noto, Japan Earthquake (EQ) 01 Jan. 2024 0.5 GSI &\& Umbra 79 8,153